### v5

# Enhancing Training Efficiency Using Packing with Flash Attention

Abstract
Padding is often used in tuning LLM models by adding special tokens to shorter training examples to match the length of the longest sequence in each batch. While this ensures uniformity for batch processing, it introduces inefficiencies by including irrelevant padding tokens in the computation and wastes GPU resources. Hugging Face trainer has always offered the option to use packing to combine multiple training examples, allowing for maximal utilization of GPU resources. However, up till now, it did not offer proper masking of each packed training example. This capability has now been added to Hugging Face Transformers 4.44. We analyse this new feature and show the benefits across different variations of packing.
1 Introduction
In LLM tasks, data is represented as sequences of tokens, where each token typically corresponds to a word, character, or sub-word. These sequences form the input to the models, which aim to learn meaningful representations and capture intricate patterns within the data. However, the lengths of these sequences can vary substantially, posing a computational challenge during training.
To illustrate this challenge, consider a scenario where we aim to fine-tune an LLM on a corpus of text. Each sentence in the corpus may have a different length, ranging from a few words to several dozen or even hundreds. Traditional approaches would require padding shorter sequences with special tokens up to the maximum sequence length. While this ensures uniformity for processing, it introduces inefficiencies by including irrelevant padding tokens in the computation, which not only wastes GPU resources but also can dilute the model’s learning signal.
Batch-level, or dynamic, padding improves this by organizing sequences within a batch in a manner to improve computational efficiency without sacrificing learning efficacy. This is achieved by dynamically padding sequences to the same length within a batch, up to the maximum example length in each batch, allowing for parallel processing across multiple sequences. By eliminating the need for fixed-length padding across batches, Batch-level padding minimizes wasted computation on padding tokens, leading to improvements in training efficiency and batch inference throughput.
Central to the implementation of batch-level padding is the concept of masking. Masking mechanisms enable neural network models to selectively ignore padded regions during computation, ensuring that they do not contribute to the model’s output or gradients. This enables the model to focus exclusively on the relevant parts of the input sequences, thereby preserving the integrity of the learning process while mitigating the effects of variable-length sequences. While batch-level padding is less wasteful than traditional padding, it is possible to improve computational efficiency further by packing small examples together, provided the masking mechanism is aware of the new example boundaries.
Packing with Position IDs entails concatenating sequences into a single tensor and applying masking to appropriate position IDs to disregard elements from other sequences during computation. Packing improves upon batch-level padding, accelerating the training process, allowing one to experiment with larger datasets and more complex models. In summary, Packing with Position IDs offers a pragmatic solution to the challenge of processing variable-length sequences efficiently and unlocks higher levels of performance and scalability to LLM sequence processing. We provide a solution for this as well as a detailed analysis of its benefits.
2 Related Work
There are a few ways that sample packing can be enabled with Flash Attention 2 [9] and proper attention masking of each example. The Flash Attention repository itself offers a way to pack while enabling proper masking of examples with Flash Attention. See Flash Attention Closed Issue 654 [1]. Others have proposed padding-free transformers, such as [6, 7]. The padding-free transformer methods require substantial and intrusive changes however to Hugging Face transformers library. As such, these methods have seen less uptake by the community and are not currently available in the Hugging Face library. For Hugging Face Trainer users, it is desirable to have a readily available solution without requiring going outside the library. This feature is now available in Transformers version 4.44.
Most works on sample packing are concerned with how to select the sequences to pack together. This problem can be formulated as a bin packing, or machine scheduling, problem from the combinatorial optimisation literature. For instance, [12] discuss one such approach for sequence selection and packing. In [8] the authors also use packing to combine multiple training examples into a single sequence, separating inputs from targets using an end-of-sequence token, and also using masking to prevent tokens from attending to others across the packed example boundaries.
The multi-pack sampler repository [5], used also by [3], employs a first-fit-decreasing heuristic for the bin packing problem to select sequences to put together on each gpu in a distributed computing setting. Since this method is the most widely used, we make use of it in conjunction with our Packing with PositionIDs solution to further enhance the packing performance.
A related effort for pre-training is studied by [pack-pretrain] who note that using causal loss masking across different documents degrades performance of the models. The authors propose to limit causal masking to within documents during pre-training. We mention also the LengthGroupedSampler function [4] in Hugging Face Transformers library which is often used in conjunction with sample padding. We consider this method as another baseline, referred to as GroupByLength+Padding.
3 Packing with Position IDs
We denote the input_ids of the tokenized $i$-th example as a tensor of shape $(L_{i},\,)$:
$\mathbf{X}_{i}=[\mathbf{x}_{i}^{(0)},\cdots,\mathbf{x}_{i}^{(L_{i}-1)}].$
Let $\mathbf{X}_{i_{1}},\mathbf{X}_{i_{2}},\cdots$ be an ordering of the training examples. Assuming a batch size $bs=4$, consider a batch consisting of examples $\{\mathbf{X}_{i_{1}},\mathbf{X}_{i_{2}},\mathbf{X}_{i_{3}},\mathbf{X}_{i_{4}}\}$. Then, in a batch-level padding based approach, the batch, $B$, is processed as a tensor of $(4,L_{max})$, where $L_{max}:=max\{L_{i_{1}},L_{i_{2}},L_{i_{3}},L_{i_{4}}\}$ and the required number of padding tokens are appended to each example ($L_{max}-L_{i_{1}}$ to $\mathbf{X}_{i_{1}}$) to make them tensors with shape $(L_{max},)$.
To avoid cross-contamination while packing examples, we propose to utilize position IDs to demarcate boundaries of individual examples in a packed sequence. We assume support for Flash Attention and the availability of position IDs which is the case for the most popular open-source LLM models.
Then, to use the position IDs, we require arranging the data sequences accordingly. There are multiple ways to arrange the data with this solution, we discuss three below.
3.1 Online minibatch Collating
In minibatch collating, the padding-free collator must pack the examples online, for each minibatch, into a tensor of dimension, $\dim_{mini}$:
$\dim_{mini}=(1,\sum_{i=1\ldots 4}L_{i}),$
and then provide position IDs which then take the form:
[…]
This capability has been added to Hugging Face Trainer from Transformers 4.44, leveraging the new DataCollatorWithFlattening. Hugging Face users of the TRL library can also benefit from this capability by setting a new flag padding_free=true in the DataCollatorForCompletionOnlyLM function.
3.2 Offline Batch Collating
It is also possible to pack a full set of samples offline in a single tensor of dimension $\dim_{flat}$:
$\dim_{flat}=(1,bs*msl),$
where $bs$ is the batch size and $msl$ denotes the maximum sequence length allowed for individual training examples.
3.3 Collating with Optimised Sample Selection
One may wish to leverage bin-packing-type sample selection algorithms such as [5] in conjunction with the solution provided by Packing with PositionIDs. In this case the samples to assign to each gpu and the position IDs for those samples are grouped per gpu. We call this Multipack+Position IDs. To contrast with the Multipack bin packing method, we introduce two simple baselines: (q) RandomPacking+PosID and (b) SortedPacking+PosID; Examples are packed on a first-come-first-serve basis from randomly ordered training data set and sorted (length-wise from long to short) data set, respectively.
Method Attention Using Correct Cross- No Broken No Sorting
Position IDs Attention Examples of Dataset
RandomSampling+Padding
GroupByLength+Padding
MiniBatchPacking+PosIDd
FixedLengthPacking
FixedLengthPacking+PosID
Multipack+PosID
SortedPacking+PosID
RandomPacking+PosID
Table 1: Characteristics of different methods of grouping examples into batches
Specifically, to enable the use of position IDs, we modify the models’ _flash_attention_forward(), adding the argument position_ids and extracting the number of examples in the batch from the position_ids. When attention_mask is None in the case of number of examples > batch size, we compute cu_seq_len from position_ids and use the flash_attn_varlen_func().
3.4 PaddingFreeCollator
We provide a new off-the-shelf data collator, the PaddingFreeCollator, summarised below.
return { input_ids: …,  labels: …,  position_ids: … }
input_ids (1, $\sum$ sequence_length_of_each_example) :
Concatenate all the examples in the mini.
labels (1, $\sum$ sequence_length_of_each_examples) :
Convert the first label_id of each example into -100, then concatenate.
position_ids (1, $\sum$ sequence_length_of_each_examples) :
Generate position ids for each example and concatenate them all.
Next we illustrate the simplicity of the method with an example, below.
3.5 Example
Consider again the scenario with a batchsize = 4 where the sequences are as follows:
[[10,11,12,13],[[10,11,12,13],[ [ 10 , 11 , 12 , 13 ] ,
$[20,21,22,23,24,25,26,27],$
$[30,31,32,33,34],$
[40,41,42,43,44,45,46,47,48,49,410]][40,41,42,43,44,45,46,47,48,49,410]][ 40 , 41 , 42 , 43 , 44 , 45 , 46 , 47 , 48 , 49 , 410 ] ]
The padding-free collator returns the input IDs, labels, and the position IDs of each example after concatenating the examples together. Hence, the collator provides:
input_ids:
[[10,11,12,13,20,21,22,23,24,25,26,27,[[10,11,12,13,20,21,22,23,24,25,26,27,[ [ 10 , 11 , 12 , 13 , 20 , 21 , 22 , 23 , 24 , 25 , 26 , 27 ,
30,31,32,33,34,40,41,42,43,44,45,46,47,48,49,410]],30,31,32,33,34,40,41,42,43,44,45,46,47,48,49,410]],30 , 31 , 32 , 33 , 34 , 40 , 41 , 42 , 43 , 44 , 45 , 46 , 47 , 48 , 49 , 410 ] ] ,
labels:
[[100,11,12,13,100,21,22,23,24,25,26,27,[[-100,11,12,13,-100,21,22,23,24,25,26,27,[ [ - 100 , 11 , 12 , 13 , - 100 , 21 , 22 , 23 , 24 , 25 , 26 , 27 ,
100,31,32,33,34,100,41,42,43,44,45,46,47,48,49,410]],-100,31,32,33,34,-100,41,42,43,44,45,46,47,48,49,410]],- 100 , 31 , 32 , 33 , 34 , - 100 , 41 , 42 , 43 , 44 , 45 , 46 , 47 , 48 , 49 , 410 ] ] ,
position_ids:
[[0,1,2,3,0,1,2,3,4,5,6,7,[[0,1,2,3,0,1,2,3,4,5,6,7,[ [ 0 , 1 , 2 , 3 , 0 , 1 , 2 , 3 , 4 , 5 , 6 , 7 ,
0,1,2,3,4,0,1,2,3,4,5,6,7,8,9,10]].0,1,2,3,4,0,1,2,3,4,5,6,7,8,9,10]].0 , 1 , 2 , 3 , 4 , 0 , 1 , 2 , 3 , 4 , 5 , 6 , 7 , 8 , 9 , 10 ] ] .
The modifications required are lightweight and are limited to providing the position IDs to Flash Attention. For details, see the repository at [2]. This solution relies, however, on the model exposing position IDs. As of the time of writing, 14 models expose position IDs and are supported by the solution. Specifically: llama, mistral, mixtral, granite, dbrx, falcon, gemma, olmo, phi, phi2, phi3, qwen2, qwen2_moe, stablelm, starcoder2 are all supported by the solution.
4 Experiments and Results
We consider three different datasets with different example-length characteristics: FLAN [15], OrcaMath[13], and python code data from The Stack[11]. A subset of 20K examples from each of the 3 datasets was used in the experiments. We evaluate the throughput in tokens/second, denoted Tok/s, peak memory, denoted Mem, and validation loss, denoted VLoss. All tests were performed on a single A100-80GB node having 8 GPU with FSDP. The throughput measurements are averaged over the 8 GPU and the peak memory recorded is the maximum over the 8 GPU.
Figure 1: Histograms of sequence lengths for the 3 training datasets: (top left) FLAN_20k, (top right) OrcaMath_20k, and (bottom) the Stack_20k.
The characteristics of example lengths in our 20K subsets of FLAN, OrcaMath, as well as in the python coding dataset from The Stack are shown in the histograms in Figure 1. Observe that both FLAN and OrcaMath have primarily small example lengths, which makes them amenable to significant throughput gains via packing, while the coding dataset the Stack has significantly longer examples, and hence we expect less throughput benefit from packing on that dataset.
4.1 Main Results
Four approaches were evaluated. The two base cases are: (i) No packing with truncate, i.e., padding, and (ii) Basic packing (as is currently available in SFT trainer, without position IDs). Then, we evaluate two variants of the proposed solution: (iii) Our solution providing packing with position IDs, in which the batch is prepared offline and flattened, and (iv) Our solution with online packing with position IDs for each minibatch.
We expect that base case (i) using padding will be the slowest approach but provide optimal training loss reduction since it does not disrupt any training examples. On the other hand, basic packing without position IDs can achieve high throughput, at the expense of distorting examples through improper attention masks.
We demonstrate the benefits of the solution on 10 of the 14 supported models. Specifically, we test it on Llama-2-7B-fp16, mistralai/Mistral-7B-v0.1, granite-8b-code-base, tiiuae/falcon-7b, google/gemma-7b, microsoft/phi-2, Qwen/CodeQwen1.5-7B, bigcode/starcoder2-7b, stabilityai/stablelm-2-1_6b, Qwen/Qwen1.5-MoE-A2.7B.
Table 2 provides the main results on a 20K subset of the FLAN instruct tuning dataset [15] and Table 3 shows the analogous results on the OrcaMath dataset [13]. For both datasets, all models are run for one epoch with gradient accumulation steps (gas) of 2, maximum sequence length (msl) of 4096 and minibatch size (bs) per GPU set to 4.
Observe that the Packing with PositionID methods improve the throughput substantially in terms of tokens per second, in many case above and beyond even basic packing that does not account for improper cross-attention (the second row in each model section).
However, this maximal packing has an impact on the loss behaviour. Due to the fact that far fewer optimisation steps are taken with such maximal packing, the loss does not decrease as fast, and its effect is confirmed by the validation loss ("VLoss") after one epoch. Hence, as a best-of-both worlds remedy, we propose the online minibatch approach to Packing with PositionIDs. Due to the packing of fewer examples in each pack, minibatch packing does not achieve the maximal throughput, however it achieves the same optimal loss pattern and hence validation loss, as the inefficient padding-based approach. It also improves on throughput substantially beyond the padding-based approach.
The variation in benefit from our proposed approach across the FLAN and OrcaMath datasets can be attributed to the statistical dataset characteristics of each; the very small sample lengths of FLAN, whose primary mode is around 100 tokens, is highly amenable to significant throughput increase from packing, more so than OrcaMath, whose mode is around 400 tokens, as seen in Figure 1. This means that offline packing can achieve very significant throughput improvements. On the other hand, with minibatch packing, the padding baseline pads only to the length of the longest sequence in each minibatch. Hence, the variance (rather than the mean) of the sample lengths affects the gain in throughput, with high variance leading to more minibatch padding. This is also confirmed by the higher improvement in throughput from minibatch packing with PositionIDs compared to padding in FLAN, Table 2 versus OrcaMath, in Table 3. The benefits across model architectures are consistent, with the exception of Gemma-7B and Qwen1.5-MoE-A2.7B.
Model Pack? inp_ids pos_id? Rows Time Tok/s Mem VLoss
llama2 no bs, $\max(B)$ no 19958 1526 771 29234 1.266
yes bs, seq no 2346 388 3091 29433 1.579
flat 1, bs*msl 1, bs*msl 599 366 3222 29438 1.578
mini 1, $\sum(B)$ 1, $\sum(B)$ 19958 809 1455 23854 1.262
mistral no bs, $\max(B)$ no 19961 1553 742 30625 1.129
yes bs, seq no 2294 392 2986 30783 1.306
flat 1, bs*msl 1, bs*msl 585 383 3010 30783 1.284
mini 1, $\sum(B)$ 1, $\sum(B)$ 19961 818 1408 24549 1.127
granite no bs, $\max(B)$ no 19947 1741 693 36277 1.169
yes bs, seq no 2420 471 2628 36278 1.740
flat 1, bs*msl 1, bs*msl 620 459 2657 36276 1.767
mini 1, $\sum(B)$ 1, $\sum(B)$ 19947 888 1358 27538 1.169
falcon no bs, $\max(B)$ no 19969 1229 890 29894 1.842
yes bs, seq no 2168 326 3401 30106 2.585
flat 1, bs*msl 1, bs*msl 536 293 3639 30202 2.415
mini 1, $\sum(B)$ 1, $\sum(B)$ 19969 687 1592 24622 1.841
gemma no bs, $\max(B)$ no 19974 1671 602 49982 1.930
yes bs, seq no 1990 461 2211 50014 3.464
flat 1, bs*msl 1, bs*msl 1056 554 1824 50020 3.418
mini 1, $\sum(B)$ 1, $\sum(B)$ 19974 1223 823 41628 1.923
phi2 no bs, $\max(B)$ no 19974 741 1449 20228 2.948
yes bs, seq no 2123 187 5806 20293 3.375
flat 1, bs*msl 1, bs*msl 525 172 6066 20342 3.464
mini 1, $\sum(B)$ 1, $\sum(B)$ 19974 415 2588 13315 2.940
CodeQwen no bs, $\max(B)$ no 19966 1477 732 44549 1.578
yes bs, seq no 2151 368 2992 44575 1.895
flat 1, bs*msl 1, bs*msl 544 367 2942 44579 1.855
mini 1, $\sum(B)$ 1, $\sum(B)$ 19966 784 1380 31652 1.577
starcoder no bs, $\max(B)$ no 19947 1585 761 34122 2.085
yes bs, seq no 2420 417 2965 34122 2.547
flat 1, bs*msl 1, bs*msl 620 396 3072 34121 2.483
mini 1, $\sum(B)$ 1, $\sum(B)$ 19947 835 1445 25484 2.090
stablelm no bs, $\max(B)$ no 19969 369 2820 16488 1.064
yes bs, seq no 2064 91 11595 16682 1.252
flat 1, bs*msl 1, bs*msl 522 85 12100 16700 1.201
mini 1, $\sum(B)$ 1, $\sum(B)$ 19969 229 4540 11011 1.066
QwenMOE no bs, $\max(B)$ no 19969 4066 254 55436 1.268
yes bs, seq no 2045 497 2101 55733 1.592
flat 1, bs*msl 1, bs*msl 1083 607 1706 55691 1.556
mini 1, $\sum(B)$ 1, $\sum(B)$ 19969 4014 257 50774 1.268
Table 2: Summary of main results on a 20K sample of the FLAN dataset over 10 of the 14 supported models. The table compares padding each example (first row, each section) with basic packing without position IDs (second row, each section), and two variants of the solution with position IDs: offline flattening and the online minibatch.
Model Pack? inp_ids pos_id? Rows Time Tok/s Mem VLoss
llama2 no bs, $\max(B)$ no 20000 790 1269 22305 0.458
yes bs, msl no 1946 315 3155 29433 0.550
flat 1, bs*msl 1, bs*msl 488 295 3324 29432 0.521
mini 1, $\sum(B)$ 1, $\sum(B)$ 20000 574 1746 20950 0.459
mistral no bs, $\max(B)$ no 20000 812 1216 23603 0.321
yes bs, msl no 1921 326 3006 30783 0.353
flat 1, bs*msl 1, bs*msl 478 307 3137 30783 0.350
mini 1, $\sum(B)$ 1, $\sum(B)$ 20000 596 1658 22409 0.320
granite no bs, $\max(B)$ no 20000 809 1199 26489 0.3242
yes bs, seq no 1887 369 2614 36278 0.3928
flat 1, bs*msl 1, bs*msl 473 351 2712 36277 0.383
mini 1, $\sum(B)$ 1, $\sum(B)$ 20000 575 1689 24423 0.3242
falcon no bs, $\max(B)$ no 20000 661 1428 22770 0.618
yes bs, msl no 1834 279 3358 30106 0.744
flat 1, bs*msl 1, bs*msl 456 248 3700 30201 0.695
mini 1, $\sum(B)$ 1, $\sum(B)$ 20000 500 1887 21702 0.618
gemma no bs, $\max(B)$ no 20000 1076 858 32325 0.344
yes bs, msl no 1794 420 2188 50014 0.962
flat 1, bs*msl 1, bs*msl 919 470 1944 50020 0.798
mini 1, $\sum(B)$ 1, $\sum(B)$ 20000 949 972 30794 0.345
phi2 no bs, $\max(B)$ no 20000 371 2240 11209 0.479
yes bs, msl no 1614 143 5755 20293 0.561
flat 1, bs*msl 1, bs*msl 399 129 6225 20341 0.514
mini 1, $\sum(B)$ 1, $\sum(B)$ 20000 306 2714 9817 0.483
CodeQwen no bs, $\max(B)$ no 20000 750 1187 28760 0.483
yes bs, msl no 1730 295 3001 44575 0.541
flat 1, bs*msl 1, bs*msl 432 288 3028 44579 0.522
mini 1, $\sum(B)$ 1, $\sum(B)$ 20000 559 1594 26132 0.483
starcoder no bs, $\max(B)$ no 20000 782 1271 24475 0.504
yes bs, msl no 1933 330 2988 34122 0.563
flat 1, bs*msl 1, bs*msl 483 302 3219 34123 0.558
mini 1, $\sum(B)$ 1, $\sum(B)$ 20000 575 1729 22439 0.504
stablelm no bs, $\max(B)$ no 20000 226 3970 8986 0.534
yes bs, msl no 1744 77 11578 16682 0.573
flat 1, bs*msl 1, bs*msl 434 70 12470 16700 0.567
mini 1, $\sum(B)$ 1, $\sum(B)$ 20000 195 4614 7368 0.534
QwenMOE no bs, $\max(B)$ no 20000 4165 216 45152 0.335
yes bs, msl no 1744 444 2005 55733 0.366
flat 1, bs*msl 1, bs*msl 889 484 1826 55692 0.352
mini 1, $\sum(B)$ 1, $\sum(B)$ 20000 3849 234 44572 0.335
Table 3: Summary of main results on a 20K sample of the OrcaMath dataset over 10 of the 14 supported models. The table compares padding each example (first row, each section) with basic packing without position IDs (second row, each section), and two variants of the solution with position IDs: offline flattening and the online minibatch.
4.2 Ablation Study on Minibatch Size
In this ablation study we compare the effect of minibatch size (bs) on training throughput and memory for two approaches: the standard padding vs proposed minibatch packing. For this experiment, we fine-tune Mistral-7B[10] on 2 different training datasets: FLAN_20k, (right) OrcaMath_20k. We present the throughput results in Figure 3 and the memory results in 2. The advantage of minibatch packing is significant, not only is the throughput a factor of 2x faster and the peak memory usage significantly lower than using padding, but the training runs out of memory (OOM) much later allowing for larger batch sizes.
Figure 2: Minibatch packing offers lower GPU peak memory usage than padding approach as minibatch size increases. “OOM" denotes out of memory error. Lower is better.
Figure 3: Minibatch packing offers better throughput (Tokens/s) than padding as minibatch size increases. “OOM" denotes out of memory error. Higher is better.
4.3 Ablation Study on Sample Selection Method
In this ablation study we consider two different models: Mistral-7B[10] and CodeLlama-Python-7B [14]. In Tables 4,5, and 6, we examine different batch sizes and sequence lengths on two of the most popular models, Mistral-7B and CodeLlama-Python-7B with several sample selection methods. We provide the results on FLAN, OrcaMath, and the Stack [11], respectively. In the tables, bs is the minibatch size per GPU, msl stands for maximum sequence length of an example/packed sequence, and gas denotes the number gradient accumulation steps. Note that the Stack dataset has high variance in sequence length as seen in Figure 1, and long sequence lengths.
As expected, the padding-based approaches have the lowest training throughput, and offline packing offers the highest throughput. Also as expected and as seen previously minibatch packing offers an intermediate ground with still-signicant throughput increases.
It is instructive to delve into both training and validation loss. The offline packing methods, even using position IDs fail to provide the same level of validation loss after one epoch that padding and minibatch packing with position IDs provide. As noted above, this is due primarily to the fewer number of optimisation update steps performed.
Finally, we examine how much benefit can be achieved by coupling the use of Packing with PositionIDs with the bin-packing-based sample selection of Multi-pack sampler [5]. The first observation is that the benefit, across all three datasets, is rather limited. However, while the throughput decreases very slightly when using the Multi-pack sampler, there is on average a small benefit to the loss behaviour of the models using bin-packing-based sampling.
Method (bs, msl, gas) Steps Time Tok/s Train Loss V Loss
RandomSampling+Padding (2, 4096, 32) 311 5516 1671 1.271 1.117
GroupByLength+Padding (2, 4096, 32) 311 3997 2306 1.277 1.120
MiniBatchPacking+PosIDd (2, 4096, 32) 311 4028 2289 1.273 1.123
FixedLengthPacking (2, 4096, 32) 35 2895 3184 1.372 1.294
FixedLengthPacking (1, 8192, 32) 35 3058 3014 1.362 1.352
FixedLengthPacking+PosID (2, 4096, 32) 35 2888 3191 1.378 1.221
FixedLengthPacking+PosID (1, 8192, 32) 35 2781 3314 1.363 1.221
Multipack+PosID (1, 8192, 32) 37 2782 3314 1.335 1.215
SortedPacking+PosID (1, 8192, 32) 37 2842 3243 1.318 1.278
RandomPacking+PosID (1, 8192, 32) 37 2852 3233 1.335 1.208
FixedLengthPacking (2, 4096, 4) 281 2959 3115 1.339 1.252
FixedLengthPacking (1, 8192, 4) 281 3143 2932 1.331 1.288
FixedLengthPacking+PosID (2, 4096, 4) 281 2845 3239 1.338 1.170
FixedLengthPacking+PosID (1, 8192, 4) 281 2839 3246 1.321 1.154
Multipack+PosID (1, 8192, 4) 299 2953 3122 1.294 1.153
SortedPacking+PosID (1, 8192, 4) 300 2938 3138 1.264 1.210
RandomPacking+PosID (1, 8192, 4) 300 2912 3166 1.293 1.151
RandomSampling+Padding (4, 4096, 16) 311 6184 1491 1.295 1.127
GroupByLength+Padding (4, 4096, 16) 311 3394 2716 1.315 1.131
MiniBatchPacking+PosIDd (4, 4096, 16) 311 3310 2785 1.296 1.127
FixedLengthPacking (4, 4096, 16) 35 2886 3193 1.369 1.307
FixedLengthPacking (1, 16384, 16) 35 3042 3027 1.358 1.431
FixedLengthPacking+PosID (4, 4096, 16) 35 2699 3415 1.375 1.225
FixedLengthPacking+PosID (1, 16384, 16) 35 2683 3432 1.350 1.225
Multipack+PosID (1, 16384, 16) 36 2700 3414 1.337 1.217
SortedPacking+PosID (1, 16384, 16) 36 2688 3429 1.312 1.281
RandomPacking+PosID (1, 16384, 16) 36 2722 3386 1.332 1.223
FixedLengthPacking (4, 4096, 2) 281 2878 3202 1.340 1.255
FixedLengthPacking (1, 16384, 2) 281 3109 2961 1.324 1.363
FixedLengthPacking+PosID (4, 4096, 2) 281 2759 3340 1.340 1.169
FixedLengthPacking+PosID (1, 16384, 2) 281 2833 3250 1.313 1.145
Multipack+PosID (1, 16384, 2) 290 2772 3326 1.299 1.154
SortedPacking+PosID (1, 16384, 2) 290 2822 3267 1.268 1.193
RandomPacking+PosID (1, 16384, 2) 290 2794 3299 1.294 1.156
Table 4: Finetuning Mistral-7B for 1 epoch on FLAN_20k dataset: performance comparison of different sample packing approaches. The first group in each section uses minibatches and the second and third groups use offline packing. Throughput (train tokens/second) is improved substantially using Packing with Position IDs. Loss performance is the same for padding and minibatch packing.
Method (bs, msl, gas) Steps Time Tok/s Train Loss V Loss
RandomSampling+Padding (2, 4096, 32) 312 4195 1884 0.331 0.320
GroupByLength+Padding (2, 4096, 32) 312 3680 2148 0.335 0.326
MiniBatchPacking+PosIDd (2, 4096, 32) 312 3683 2146 0.331 0.320
FixedLengthPacking (2, 4096, 32) 30 2525 3130 0.383 0.352
FixedLengthPacking (1, 8192, 32) 30 2649 2985 0.369 0.359
FixedLengthPacking+PosID (2, 4096, 32) 30 2359 3351 0.396 0.348
FixedLengthPacking+PosID (1, 8192, 32) 30 2341 3378 0.383 0.348
Multipack+PosID (1, 8192, 32) 31 2397 3297 0.369 0.347
SortedPacking+PosID (1, 8192, 32) 31 2446 3232 0.370 0.347
RandomPacking+PosID (1, 8192, 32) 31 2401 3292 0.368 0.346
FixedLengthPacking (2, 4096, 4) 241 2552 3098 0.355 0.331
FixedLengthPacking (1, 8192, 4) 241 2676 2954 0.343 0.333
FixedLengthPacking+PosID (2, 4096, 4) 241 2443 3236 0.356 0.323
FixedLengthPacking+PosID (1, 8192, 4) 241 2494 3169 0.344 0.322
Multipack+PosID (1, 8192, 4) 248 2608 3031 0.331 0.323
SortedPacking+PosID (1, 8192, 4) 248 2460 3214 0.332 0.323
RandomPacking+PosID (1, 8192, 4) 248 2471 3199 0.330 0.323
RandomSampling+Padding (4, 4096, 16) 312 4048 1953 0.330 0.320
GroupByLength+Padding (4, 4096, 16) 312 2929 2699 0.335 0.325
MiniBatchPacking+PosIDd (4, 4096, 16) 312 3142 2516 0.330 0.320
FixedLengthPacking (4, 4096, 16) 30 2500 3163 0.383 0.352
FixedLengthPacking (1, 16384, 16) 30 2682 2945 0.363 0.374
FixedLengthPacking+PosID (4, 4096, 16) 30 2365 3343 0.396 0.348
FixedLengthPacking+PosID (1, 16384, 16) 30 2358 3349 0.376 0.347
Multipack+PosID (1, 16384, 16) 30 2264 3492 0.370 0.347
SortedPacking+PosID (1, 16384, 16) 30 2253 3509 0.371 0.348
RandomPacking+PosID (1, 16384, 16) 30 2240 3529 0.369 0.347
FixedLengthPacking (4, 4096, 2) 241 2559 3089 0.355 0.330
FixedLengthPacking (1, 16384, 2) 241 2697 2929 0.337 0.346
FixedLengthPacking+PosID (4, 4096, 2) 241 2362 3346 0.356 0.323
FixedLengthPacking+PosID (1, 16384, 2) 241 2331 3388 0.337 0.322
Multipack+PosID (1, 16384, 2) 245 2469 3202 0.330 0.323
SortedPacking+PosID (1, 16384, 2) 245 2379 3323 0.333 0.327
RandomPacking+PosID (1, 16384, 2) 245 2340 3378 0.330 0.322
Table 5: Finetuning Mistral-7B for 1 epoch on OrcaMath_20k dataset: performance comparison of different sample packing approaches. The first group in each section uses minibatches and the second and third groups use offline packing. Throughput (train tokens/second) is improved substantially using Packing with Position IDs. Loss performance is the same for padding and minibatch packing.
Method (bs, msl, gas) Steps Time Tok/s Train Loss V Loss
RandomSampling+Padding (2, 4096, 32) 312 12463 2167 0.440 0.505
GroupByLength+Padding (2, 4096, 32) 312 8886 3040 0.535 0.505
MiniBatchPacking+PosIDd (2, 4096, 32) 312 8845 3054 0.440 0.505
FixedLengthPacking (2, 4096, 32) 103 8157 3311 0.441 0.581
FixedLengthPacking (1, 8192, 32) 103 9117 2962 0.420 0.588
FixedLengthPacking+PosID (2, 4096, 32) 103 8157 3311 0.420 0.588
FixedLengthPacking+PosID (1, 8192, 32) 103 8129 3322 0.414 0.541
Multipack+PosID (1, 8192, 32) 121 8227 3283 0.393 0.539
SortedPacking+PosID (1, 8192, 32) 112 8224 3285 0.393 0.541
RandomPacking+PosID (1, 8192, 32) 122 8500 3178 0.392 0.538
FixedLengthPacking (2, 4096, 12) 274 8170 3306 0.436 0.574
FixedLengthPacking (1, 8192, 12) 274 8885 3040 0.415 0.583
FixedLengthPacking+PosID (2, 4096, 12) 274 8053 3354 0.432 0.516
FixedLengthPacking+PosID (1, 8192, 12) 274 8058 3352 0.411 0.515
Multipack+PosID (1, 8192, 12) 324 8617 3135 0.389 0.514
SortedPacking+PosID (1, 8192, 12) 299 8293 3257 0.389 0.518
RandomPacking+PosID (1, 8192, 12) 326 8119 3327 0.388 0.514
RandomSampling+Padding (4, 4096, 16) 312 17639 1531 0.410 0.508
GroupByLength+Padding (4, 4096, 16) 312 8621 3133 0.535 0.506
MiniBatchPacking+PosIDd (4, 4096, 16) 312 8380 3223 0.410 0.508
FixedLengthPacking (4, 4096, 16) 103 7984 3383 0.441 0.582
FixedLengthPacking (1, 16384, 16) 103 9919 2722 0.408 0.594
FixedLengthPacking+PosID (4, 4096, 16) 103 7647 3532 0.435 0.542
FixedLengthPacking+PosID (1, 16384, 16) 103 7814 3455 0.403 0.541
Multipack+PosID (1, 16384, 16) 112 7909 3415 0.392 0.539
SortedPacking+PosID (1, 16384, 16) 107 7850 3441 0.394 0.540
RandomPacking+PosID (1, 16384, 16) 112 7954 3396 0.393 0.539
FixedLengthPacking (4, 4096, 6) 274 8028 3365 0.435 0.574
FixedLengthPacking (1, 16384, 6) 274 9958 2711 0.402 0.593
FixedLengthPacking+PosID (4, 4096, 6) 274 7647 3532 0.432 0.516
FixedLengthPacking+PosID (1, 16384, 6) 274 8036 3360 0.400 0.515
Multipack+PosID (1, 16384, 6) 299 7849 3442 0.389 0.514
SortedPacking+PosID (1, 16384, 6) 286 7892 3422 0.391 0.518
RandomPacking+PosID (1, 16384, 6) 299 7986 3382 0.389 0.515
Table 6: Finetuning CodeLlama-Python-7B for 1 epoch on the Stack_20k dataset: performance comparison of different sample packing approaches. The first group in each section uses minibatches and the second and third groups use offline packing. Throughput (train tokens/second) is improved substantially using Packing with Position IDs. Loss performance is the same for padding and minibatch packing.
5 Conclusions
We provide a lightweight solution for packing with PositionIDs that has been integrated into Hugging Face Transformers library version 4.44 and we demonstrate its benefits.